OpenStudy (anonymous):

Simplify: (1+cot^2y)(cos2y + 1) The first cot is squared, the second is not. Thanks!

OpenStudy (loser66):

the second one is cot or cos?

OpenStudy (loser66):

\[(1+cot^2y)(cos2y+1)\]or \[(1+cot^2y)(cot2y+1)\]

OpenStudy (anonymous):

the second one is cos

OpenStudy (loser66):

\[1+cot^2y = \frac{1}{sin^2y}\] \[cos2y+1=2cos^2y\] ---------------------------- time them together to get \[\frac{2cos^2y}{sin^2y}=2cot^2y\]